3/2 is the solution of the equation 8x – 5 = 7.
3/2 is the solution of the equation 8x – 5 = 7.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The Given statement is True
Consider the equation, 8x–5 = 7
Transpose -5 to RHS then it becomes 5. 8x = 7 + 5
8x = 12
x = 12/8
x = 3/2

Thanks